The following table lists results for various proc
Answer #1

We know,
Q = W + ΔU

For constant pressure process,
W = Q -  ΔU
W = (5/2 - 3/2) * nR ΔT
W = nR ΔT

For Adiabatic Process, Q = 0
So,
W + ΔU = 0
ΔU = - W
ΔU = 3/2 nR ΔT

For Constant Volume process, W = 0
Q = W + ΔU
Q = ΔU
ΔU = 3/2 nR ΔT

For Isothermal Process, ΔU = 0
So,
Q = W
Q = nRTln(vF/vI)